OpenStudy (anonymous):

√27x^2y - √45xy^2 + √108x^2y - √125xy^2 a. xy√15xy b. 9x√3y - 8y√5x c. √135x^2y - √170xy^2 d. 3x√15y - 4y√5x

OpenStudy (mathteacher1729):

Is this from something which is being graded?

OpenStudy (anonymous):

Yes, but I am home schooled and don't understand this lesson.

OpenStudy (mathteacher1729):

Do you know how to "simplify a radical"? \[\begin{align}\sqrt{27}&=\sqrt{9\cdot3}\\&=\sqrt{9}\sqrt{3}\\&=\fbox{$3\sqrt{3}$}\end{align}\]

OpenStudy (mathteacher1729):

Can you do something similar to the other radical expressions?

OpenStudy (mathteacher1729):

http://www.mathtv.com/videos_by_topic This website is AMAZING. The old guy is the best. Click "Algebra, Simplifying Expressions, then with radicals". There are like 10 video examples.

OpenStudy (anonymous):

I know a little bit about simplifying radicals, I can add and subtract ones like these: '4√2 + 7√2 - 3√2"

OpenStudy (anonymous):

Thank you (:

OpenStudy (mathteacher1729):

So basically with this problem, what you wanna do is this: Simplify all the radicals and combine "like radicals" .

OpenStudy (mathteacher1729):

so go one step at a time, try it with \[\sqrt{45}\]

OpenStudy (anonymous):

√45 = (9 x 5) = 3√5

OpenStudy (mathteacher1729):

Yup! Keep going with the rest. :)